Subject: APO 3.5m users committee minutes, April 2011
From: Suzanne Hawley
Submitted: Wed, 13 Apr 2011 17:03:52 -0700 (PDT)
Message number: 1223
(previous: 1222,
next: 1224
up: Index)
APO 3.5-m Users Committee Phonecon, 4/04/11
Attending: Suzanne Hawley, Mark Klaene, Michael Strauss, Bill Ketzeback,
Scott Anderson, Remy Indebetouw, Sean Moran, Jon Holtzman, John Bally,
Al Harper, and Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
o New Mexico State (Jon Holtzman) - Jon had nothing to report.
o Washington (Scott Anderson) - Scott said that one of the UW users could
not display DIS slit-viewer images. This was within TUI, and seemed to
be a permissions issue. Nobody else on the phonecon reported similar
problems, so the problem may rest at UW. Mark said that there is a
problem report that seeks to get rid of permissions for slit-viewer and
guide camera images, which may be related to this problem. This also
could be because of the known "two IP addresses needed" feature, one
for login and the other for HTML. Bill Ketzeback agreed to look further
into this.
ACTION: Bill Ketzeback to look into and suggest fix to permissions
and/or IP issues in TUI related to slit-viewer and guide camera images.
o Colorado (John Bally) - John had nothing new to report
o Princeton (Michael Strauss) - Michael said everything is fine.
o Johns Hopkins (Sean Moran) - Sean had nothing new to report.
o Virginia (Remy Indebetouw) - Remy said that he couldn't get to the APO
weather pages or users manuals from a hotel. He could run TUI fine,
which is weird. Mark said that he will verify that all the APO pages
are "open".
o Chicago (Al Harper) - no report.
**********************************
Discussion of telescope/instruments report - Klaene:
Mark mentioned the highlights from the past few weeks. It was a hectic
month, and quarter, at APO. There was the GFP upgrade effort with
the NASA/Goddard folks at the site, a Colorado student group, closely
followed by the 2nd annual open house, and then the BOSS team meeting
in Cloudcroft. It has been dry and windy, more dry than usual and
that could lead to forest restrictions being put in place within weeks.
Nonetheless, with the generally clear skies a lot of observing has
gotten accomplished. The DIS slit-view camera will be removed soon
so that measurements can be taken; a rebuild is being planned.
The spontaneous Agile aborts seem to be much improved, and as far as we
are concerned, fixed as well as can be without undertaking a complete
replacement of the Agile computer and controller. The other instruments
are working well, and for NIC-FPS, the read noise is much better, but
the upside-down readout problem is still with us. Although the controller
often fails when the instrument is inverted while tracking and observing,
the recovery is quick. The GFP upgrade will continue in June.
Mark's detailed report for the reporting period follows:
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights
2/24//10 through 3/30/11
0) Overview
Weather has been generally good with occasional spring winds and elevated
dust levels. The site was busy much of March with the NASA Goddard folks
on site in early March working on GFP-IFS followed immediately by a large
group of students from Colorado, then the APO Open House which brought
about 88 members of the public, and ending with nearly 100 SDSS-III astronomers
in town for a BOSS collaboration meeting. In preparing for the open house
the APO display at the Sunspot Astronomy Visitor Center was completely
remodeled and updated. Gretchen Van Doren was a key organizer for many
of these events and did a magnificent job.
The forest remains dry and extreme forest fire danger is expected with access
restrictions starting soon.
1) Telescope
Virtually no telescope problems were reported during this period. The spare
direct-drive motor is now complete.
2) Instruments
DIS performed well this period. Work is planned for April to remove the
slit-viewer assembly to investigate new mounting options and replacing
the camera and optics.
SPICam operated with no new issues.
Agile spontaneous aborts work is complete for now. There may still be a
rare failure at the start of an exposure sequence, but changes to software
have all but eliminated the problems of the past. Also work has begun to
install a brake on the rotator. Previous work has eliminated the rotator
failures and the brake will eliminate the need for the pin and will hold
the rotator in place when the power is off.
TripleSpec is not reporting slit position in the headers. This is on
the list of things to fix.
The echelle has been running slightly warmer than normal so time has been
set aside in April to vacuum pump the Schmidt camera.
NIC-FPS was returned to the observatory after a very successful trip to CU.
Upon return, the read noise is now better than the 2 channel configuration.
However the problem of spontaneous reset of the Leach controller when rotated
upside still exists. Users are advised to start their observations at
specific recommended rotation angles, depending on sky position, to
avoid tracking upside down. There is also a software alert in place that
should trigger if an image is bad. See more below.
GFP upgrade started on the mountain in February. After a return trip by the
Goddard folks in March the direct imaging and IFS modes are working. The
FPS and photon-counting modes are still under development as is TUI integration.
We are hopeful that in the scheduled June engineering run they will make
further significant progress. The instrument is currently not available
for use.
**********************************
current NIC-FPS readnoise, gain, linearity information
(recommended observing is with NFS=8)
NFS readnoise
1 30.0
2 20.9
3 18.3
4 15.3
5 14.4
6 13.5
7 12.6
8 12.4
Gain = 2.94e/DN
linear to within 3% up to 54,400 raw counts and at 5% up to 56,800 counts.
------------------
NIC-FPS rotation recommendations depending on sky position of object:
Basically, the northeast is the problem quadrant; that's where the instrument
will be upside-down at default 0 object rotation.
Sky quadrant recommended object rotation
northeast +-180 (positive better if observer plans to track for awhile)
southeast to south -90
southwest 0
northwest +90
Note: the rotation needs to be changed for every target. For example, when
leaving program star and moving to standard star, users must put in new
rotation appropriate for standard star's position in the sky. Best way
to see where the next target will be is from the TUI window showing the
telescope position on the sky (window is labeled "SKY"). If in doubt,
ask the Observing Specialist!
The current plan is to take the instrument back to CU during the summer
shutdown and replace all of the detector cabling inside the dewar with
the hope that this will fix the rotation problem.
**********************************
NIC-FPS status - Hawley:
Suzanne had three things to mention:
o NIC-FPS characterization - see the report, and we're pleased with the
reduced noise performance. NIC-FPS is back in service.
o the rotation issue - should we set rotation limits in software? It is
hard to imagine simple tools for the users to avoid the inverted rotation
angles (50 to 60 degrees range around -180), although Sky View will show
you this information for the next target on your list. Observers need to
watch their images closely to see if the controller is hung up; a warning
message is now displayed. NIC-FPS users need to be especially vigilant
while observing.
ACTION: Users committee members to advise NIC-FPS users about
instrument rotation angle issue and need for vigilance.
o Stephane Beland will work on the rotation problem this summer. After
19 June, NIC-FPS will return to CU and will have all its internal cabling
replaced during July. A message to this effect will go out to the
schedulers next month.
**********************************
GFP upgrade - Klaene:
Mark reported that the GFP upgrade is in the early stages of commissioning.
The Goddard team is working to get the new features functioning through the
instrument control computer, after which some of the new capabilities will
be implemented in a TUI interface. A concentrated effort to work on
commissioning is planned for June. The new capabilities include an integral
field unit, a new detector with 3 electrons read noise, and a lenslet array.
The GFP throughput is now estimated to be 20%. There is still a fair amount
of work left to do. Suzanne added that the Users Committee members should
share this information with prospective users, and collect feedback. The
slides that were included with the agenda give more details on FOV, etc. and
can be shared by committee members with interested users. Michael suggested
that the new GFP detector could be used for direct imaging; even though the
field is relatively small, the lower read noise might make it advantageous
for some projects. Suzanne said if anyone has interest in using the upgraded
GFP for early science in Q3, they should let her know and she will make sure
contact is made with the Goddard group. The earliest return of GFP for regular
science use will likely be in Q4.
**********************************
Summer shutdown schedule - Klaene, Hawley:
Mark said that planning for the summer shutdown is progressing, and will
include a realumization of the primary mirror at Kitt Peak. The shutdown
will begin on 5 July and nominally end on 28 July. The last five days of
the shutdown involve on-sky testing before returning to routine science
observing. Given the typical weather vagaries in late July, the actual
return-to-science date could slip.
**********************************
Update on logging scripts - Ketzeback:
Mark noted that the logging scripts for DIS are in good shape, and are being
considered for some of the other instruments. Bill said that the DIS scripts
match science images with the corresponding guider images. Adam Kowalski (grad
student at UW) wrote most of the core code for the DIS script, and Bill did
some additional work on it. Bill provided a description of the DIS script
[see note below]. John Bally, who was instrumental in requesting this capability,
said that he's looking forward to using it. Suzanne mentioned that we still
need to understand what is required to implement the logging script working for
TripleSpec. The echelle would be another possible instrument to consider.
Suzanne added that we would like to get user feedback on the logging scripts.
ACTION: Remy Indebetouw to get the logging script specifications for TripleSpec.
[note added, from Bill Ketzeback]:
Adam Kowalski and Bill Ketzeback have a program running on newton.apo.nmsu.edu
for matching up DIS exposures with corresponding slit viewer images. To run
the program, log into your institutional account on newton and in the home
directory (or subdirectory) type: "dcam_match". The program will prompt for
the program id (example: Q1PU04) and the ut date of the observations (in
the form YYMMDD, example 110327). The program will output a unique log file
with information in the title (example: dcam_match_110327_Q1PU04.out).
The contents of the file contains a header explaining each of the columns.
The following is an example output for a recent program.
# UT110327, number of science files = 11
# Column 1: Science file (only IMAGETYP=object are included)
# Column 2: Nearest slitviewer image
# Column 3: MJD (hours) for the midtime of the science file
# Column 4: Midtime of science image - midtime of nearest slitviewer image
(seconds)
# Column 5: Range of slitviewer images that have mid-times within the
exposure of the science image
dvb.0001r.fits d0058.fits 7.8607 15.22 d0042.fits - d0075.fits
dvb.0002r.fits d0094.fits 8.2656 9.94 d0086.fits - d0102.fits
dvb.0003r.fits d0121.fits 8.4737 -0.56 d0120.fits - d0122.fits
dvb.0004r.fits d0146.fits 8.7376 -14.58 d0129.fits - d0162.fits
dvb.0005r.fits d0186.fits 9.2046 -0.71 d0170.fits - d0202.fits
dvb.0006r.fits d0215.fits 9.5523 14.51 d0207.fits - d0223.fits
dvb.0007r.fits d0247.fits 10.0089 9.18 d0230.fits - d0264.fits
dvb.0008r.fits d0302.fits 10.5684 -17.61 d0293.fits - d0310.fits
dvb.0009r.fits d0322.fits 10.8101 13.25 d0314.fits - d0330.fits
dvb.0010r.fits d0340.fits 11.1214 -36.44 None
dvb.0011r.fits d0341.fits 11.1598 54.43 None
For more experienced users, the user id and ut date can be entered on the
command line, for example:
dcam_match Q1PU04 110327
Plans for similar scripts for the other instrument-guider pairs are in the works.
**********************************
Echelle upgrade - Holtzman:
With help from the instrument development group at JHU, Jon hopes to identify,
acquire, and install a more efficient grating for the echelle this summer.
Other upgrade options including a new detector are on hold at the moment,
pending additional effort to determine if the payback will be worth the
cost/risk.
**********************************
Miscellaneous - Hawley:
Suzanne gave her appreciation to the site staff, Gretchen Van Doren in
particular, for all the support provided for recent site activities.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: User committee members to poll their users to
determine if astronomers and/or students are willing to help with
the ARCSAT upgrade proposal and/or its implementation if funded, and
for what scientific objectives?
There was no feedback from Princeton, Colorado or Virginia. NMSU is
interested, as are several people from UW. There is some interest at
Chicago, and Sean will check again at JHU. Suzanne said that we need
"interest" that includes a strong motivation to get involved; we'll keep
this action open for another month.
[new actions from this meeting]:
===> ACTION: Bill Ketzeback to look into and suggest fix to
permissions and/or IP issues in TUI related to slit-viewer and guide
camera images.
===> ACTION: Users committee members to advise NIC-FPS users
about instrument rotation angle issue and need for vigilance.
===> ACTION: Remy Indebetouw to get the logging script
specifications for TripleSpec.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 9 May at
8:30 AM Pacific Time. The agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1223 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO